### v4

# Bayesian Ego-graph Inference for Networked Multi-Agent Reinforcement Learning

Abstract
In networked multi-agent reinforcement learning (Networked-MARL), decentralized agents must act autonomously under local observability and constrained communication over fixed physical graphs. Existing methods often assume static neighborhoods, limiting adaptability to dynamic or heterogeneous environments. While centralized frameworks can learn dynamic graphs, their reliance on global state access and centralized infrastructure is impractical in real-world decentralized systems. We propose a stochastic graph-based policy for Networked-MARL, where each agent conditions its decision on a sampled subgraph over its local physical neighborhood. Building on this formulation, we introduce BayesG, a decentralized actor–critic framework that learns sparse, context-aware interaction structures via Bayesian variational inference. Each agent operates over an ego-graph and samples a latent communication mask to guide message passing and policy computation. The variational distribution is trained end-to-end alongside the policy using an evidence lower bound (ELBO) objective, enabling agents to jointly learn both interaction topology and decision-making strategies. BayesG outperforms strong MARL baselines on large-scale traffic control tasks with up to 167 agents, demonstrating superior scalability, efficiency, and performance.
1 Introduction
Multi-agent reinforcement learning (MARL) has emerged as a powerful framework for sequential decision-making in distributed systems, enabling applications in autonomous driving [1, 2], wireless communication [3, 4], multiplayer games [5, 6], and urban traffic control [7, 8]. A popular training paradigm is centralized training with decentralized execution (CTDE), where a centralized critic leverages global state information to train decentralized  [9, 10, 11, 12].
While CTDE methods demonstrate strong empirical performance in simulation benchmarks [13, 14, 15], they often rely on access to global observations and centralized learning infrastructure, as shown in Figure 1(a). This assumption rarely holds in real-world applications, where agents are geographically distributed and subject to local sensing and communication constraints. Instead, many practical domains—from urban mobility to smart grids—are better modeled as networked MARL [16, 17, 18, 19], where agents interact over a fixed communication graph and can only observe or exchange information with nearby neighbors.
A major challenge in networked MARL is the use of static communication graphs, where agents are hardwired to exchange information with all local neighbours regardless of contextual relevance [16, 17]. This can lead to inefficient coordination, unnecessary message exchange, or even performance degradation, particularly in dynamic settings like traffic control, where congestion varies over time and not all neighbours are equally informative. This raises a fundamental question:
Can decentralized agents learn to dynamically adapt their interaction structure using local observations and task feedback?
While recent CTDE methods have explored learning dynamic interaction graphs [20, 15, 21, 22], their applicability is limited in decentralized settings, where agents must reason over local observability and adhere to the physical structure of the environment.
To address this, we propose BayesG, a decentralized graph-based actor–critic framework for networked MARL that learns latent interaction structures via Bayesian inference. As illustrated in Figure 1(b), we begin by introducing a graph-based policy, where each agent conditions its decisions on a stochastic subgraph sampled from a learned distribution over its physical neighbourhood. Each agent operates over an ego-graph, a localized subgraph capturing its immediate neighborhood, enabling context-aware decisions under topological constraints. We formulate this process as Bayesian variational inference of latent graphs, where each agent $i$ infers a binary mask $Z_{i}$ over its local neighbourhood. The mask is treated as a latent variable with posterior $p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})$, conditioned on the physical subgraph $G^{\text{env}}_{\mathcal{V}_{i}}$ and agent-specific data $D_{i}$ (e.g., neighbour states, trajectories and polices). The posterior is approximated by a variational distribution $q(Z_{i};\phi_{i})$, optimized end-to-end via the evidence lower bound (ELBO).
BayesG integrates latent graph inference into policy learning, enabling agents to prioritize critical communication links within their local ego-graphs and prune irrelevant ones—all without requiring global supervision. This leads to task-adaptive, uncertainty-aware, and communication-efficient coordination under topological constraints. Experiments on both synthetic and real-world traffic control benchmarks show that BayesG outperforms state-of-the-art MARL baselines in both performance and interpretability.
Our main contributions are:
  • We propose a stochastic graph-based policy for networked MARL, where each agent conditions decisions on a sampled subgraph over its physical neighbourhood.
  • We formulate latent graph learning as Bayesian variational inference, treating edge masks as posterior distributions constrained by the environment topology and agent-local data.
  • We develop an end-to-end training algorithm that integrates variational graph inference with actor–critic learning via an ELBO objective.
Figure 1: (a) In CTDE, the global state is available for learning both the centralized critic and the interaction graph. (b) Overview of BayesG. In networked MARL, each agent’s state and action are influenced by its neighbors, forming local data $D_{i}=\{s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}}\}$. We formulate latent graph learning as Bayesian variational inference, where each agent infers a binary mask $Z_{i}$ over its neighborhood from the environment graph $G^{\text{env}}_{\mathcal{V}_{i}}$ and local data $D_{i}$. The posterior $p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})$ is approximated by a variational distribution $q(Z_{i};\phi_{i})$, from which a sparse subgraph is sampled and used for graph-conditioned policy learning via an ELBO objective.
2 Related work
Networked MARL. Networked MARL focuses on decentralized learning over fixed topologies, often without explicitly adapting the communication structure. Consensus-based methods [16, 17] synchronize local value functions via neighborhood averaging, typically assuming partial or global observability. Other works incorporate network-aware priors such as distance-based decay [23] or local reward aggregation [19] to promote spatially coherent coordination. Model-based frameworks [24, 25] exchange predicted trajectories but still rely on static interaction graphs. Communication-based methods like NeurComm [7] enable information sharing across neighbors, yet use fixed topologies throughout training. Recent work [26, 23, 27] has explored sampling strategies to improve scalability in networked systems, but these methods sample from predefined or uniform distributions at the environment or agent level. Our method differs fundamentally by introducing a Bayesian latent graph inference mechanism that allows each agent to adaptively select task-relevant neighbors from its physical graph, enabling communication-efficient, locally grounded coordination.
Graph-based MARL and Latent Structure Learning. Recent advances in cooperative MARL leverage latent interaction graphs to enhance coordination and scalability. Methods such as DGN [28], DICG [29], and HGAP [30] employ attention mechanisms or graph neural networks (GNNs) [31, 32] to infer inter-agent dependencies and guide information exchange. Variational approaches [20, 22] further treat interaction graphs as latent variables, jointly optimizing communication structure and policy learning. However, most existing methods, including Dec-POMDP-based coordination graph approaches [33, 29, 34, 14, 21], assume centralized training with access to global state and reward signals, and allow unrestricted graph rewiring, which are impractical in real-world decentralized settings. In contrast, we address the more realistic networked MARL scenario by learning dynamic latent interaction masks on the environment graph using only local observations.
3 Preliminaries
We consider networked MARL, over a fixed interaction graph $G=(\mathcal{V},\mathcal{E})$, where each agent $i\in\mathcal{V}$ interacts locally with neighbors $\mathcal{N}_{i}=\{j\mid(i,j)\in\mathcal{E}\}$ and defines a closed neighborhood $\mathcal{V}_{i}:=\mathcal{N}_{i}\cup\{i\}$. At each step, agent $i$ receives a local observation $\tilde{s}_{i}=f(s_{\mathcal{V}_{i}})$, which may include its own state and aggregated information from neighbors, and selects an action $u_{i}\in\mathcal{U}^{i}$ according to a local policy $\pi_{i}(u_{i}\mid\tilde{s}_{i})$. The encoder $f(\cdot)$ captures neighbourhood-level context using neural architectures (e.g., MLPs [7]); specific choices are detailed in Section 5.1.2. We begin with the spatiotemporal MDP formulation, followed by the decentralized actor–critic (A2C) training objective.
Definition 1 (Spatiotemporal MDP[7]).
A spatiotemporal MDP is defined as the tuple $(\mathcal{S},\{\mathcal{U}^{i}\}_{i\in\mathcal{V}},\mathcal{P},\mathcal{R},\gamma,\zeta)$, where $\mathcal{S}$ is the global state space, $\mathcal{U}^{i}$ the action space of agent $i$, and $\zeta$ the initial state distribution. The local transition model of agent $i$ is:
$p_{i}(s_{i}^{\prime}\mid s_{\mathcal{V}_{i}},u_{i})=\sum_{u_{\mathcal{N}_{i}}}\left(\prod_{j\in\mathcal{N}_{i}}\pi_{j}(u_{j}\mid f(s_{\mathcal{V}_{j}}))\right)p(s_{i}^{\prime}\mid s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}}),$ (1)
where the transition depends on local neighborhood states and actions. The local reward is $\mathcal{R}(s_{\mathcal{V}_{i}},u_{\mathcal{V}_{i}})\in\mathbb{R}$, and the objective is to maximize expected return: $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{\mathcal{V}_{i},t},u_{\mathcal{V}_{i},t})\right]$, with discount factor $\gamma\in(0,1]$.
Definition 2 (Networked MARL with decentralized A2C[7]).
Let $\{\pi_{\theta_{i}}\}_{i\in\mathcal{V}}$ and $\{V_{\omega_{i}}\}_{i\in\mathcal{V}}$ be decentralized actor and critic networks, respectively. Given an on-policy minibatch $\mathcal{B}=\{(s_{i,\tau},u_{i,\tau},r_{i,\tau},u_{\mathcal{N}_{i},\tau})\}$, the losses are:
$\mathcal{L}(\theta_{i})$ $=\frac{1}{|\mathcal{B}|}\sum_{\tau\in\mathcal{B}}\left[-\log\pi_{\theta_{i}}(u_{i,\tau}\mid\tilde{s}_{i,\tau})\hat{A}_{i,\tau}^{\pi}+\beta\sum_{u_{i}\in\mathcal{U}^{i}}\pi_{\theta_{i}}(u_{i}\mid\tilde{s}_{i,\tau})\log\pi_{\theta_{i}}(u_{i}\mid\tilde{s}_{i,\tau})\right],$ (2)
$\mathcal{L}(\omega_{i})$ $=\frac{1}{|\mathcal{B}|}\sum_{\tau\in\mathcal{B}}\left(\hat{R}_{i,\tau}^{\pi}-V_{\omega_{i}}(\tilde{s}_{i,\tau},u_{\mathcal{N}_{i},\tau})\right)^{2},$ (3)
where $\tilde{s}_{i,\tau}=f(s_{\mathcal{V}_{i},\tau})$ is the encoded local observation. The $\hat{A}_{i,\tau}^{\pi}=\hat{R}_{i,\tau}^{\pi}-v_{i,\tau}$ is advantage estimate, where the reward is $\hat{R}_{i,\tau}^{\pi}=\sum_{\kappa=0}^{K-1}\gamma^{\kappa}\sum_{j\in\mathcal{V}_{i}}\alpha^{d_{ij}}r_{j,\tau+\kappa}+\gamma^{K}v_{i,\tau+K},$, and $K$ denotes the rollout horizon. The $v_{i,\tau}=V_{\omega_{i}^{\prime}}(\tilde{s}_{i,\tau},u_{\mathcal{N}_{i},\tau})$ is the target critic output. The $\alpha\in(0,1]$ adjusts influence from distant neighbors, and $\beta$ controls the entropy regularization.
4 Method
We propose BayesG, a decentralized actor–critic framework for networked MARL that learns sparse, context-aware interaction graphs via variational inference. We begin by formulating a graph-based policy, where each agent conditions its decision on a sampled subgraph over its physical neighborhood. This formulation enables agents to adaptively modulate their coordination structure based on local observations while respecting communication constraints. We then extend this framework to optimize both the policy and graph sampling distribution jointly using a variational learning objective.
4.1 Graph-based Policy with Latent Interaction Structures
In conventional A2C for networked MARL (Definition 2), agents condition their policy on local observations $\tilde{s}_{i,\tau}=f(s_{\mathcal{V}_{i},\tau})$, which implicitly encode neighbor influence but treat all neighbors as equally informative. This uniform treatment limits the agent’s ability to adapt to dynamic environments or prioritize critical interactions. To overcome this, we introduce a graph-based policy in which each agent samples a subgraph from a learned distribution and conditions its decision on the sampled structure. (See Appendix A.1 for details).
Definition 3 (Graph-based Policy).
The policy of agent $i$ is defined as:
$\pi_{i}(u_{i},G_{\mathcal{V}_{i}}\mid s_{\mathcal{V}_{i}};\theta_{i},\varphi_{i})=\rho(G_{\mathcal{V}_{i}}\mid s_{\mathcal{V}_{i}};\varphi_{i})\cdot\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}});\theta_{i}),$ (4)
where $G_{\mathcal{V}_{i}}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$ is a sampled binary adjacency matrix drawn from distribution $\rho(\cdot;\varphi_{i})$ over agent $i$’s closed neighborhood. The graph is restricted to the environment topology: $G_{\mathcal{V}_{i}}\in\mathcal{G}^{\text{env}}_{\mathcal{V}_{i}}:=\{G\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}\mid G\preceq G^{\mathrm{env}}_{\mathcal{V}_{i}}\}$, ensuring only physically permitted connections. The function $\tilde{f}_{i}(\cdot)$ is a graph-conditioned encoder, and $\tilde{\pi}_{i}(\cdot\mid\cdot;\theta_{i})$ is the action-selection policy.
This formulation enables agents to stochastically select contextually relevant neighbors, driven by local observations. By formulating coordination as a distribution $\rho(\cdot)$ over subgraphs, this approach provides a general framework for inferring context-dependent interaction structures—allowing agents to explore diverse coordination patterns and adapt to non-stationary dynamics. In traffic signal control, this allows dynamically emphasizing congested junctions while suppressing irrelevant neighbors.
We now extend the A2C training objective to accommodate graph-conditioned policies. This graph-based loss serves as the actor component of our training framework and jointly optimizes the policy and interaction structure. (See Appendix A.2 for details).
Definition 4 (Graph-based A2C Objective).
$\mathcal{L}_{\theta,\varphi}=\frac{1}{|\mathcal{B}|}\sum_{\tau\in\mathcal{B}}\mathbb{E}_{G_{\mathcal{V}_{i}}\sim\rho}\left[-\log\tilde{\pi}_{i}(a_{i,\tau}\mid\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}}))\cdot\hat{A}_{i,\tau}^{\pi}+\beta\cdot\mathcal{H}(\tilde{\pi}_{i}(\cdot\mid\tilde{f}_{i}))\right],$ (5)
where the entropy term is defined as $\mathcal{H}(\tilde{\pi}_{i}(\cdot\mid\tilde{f}_{i}))=-\sum_{u_{i}\in\mathcal{U}^{i}}\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i})\log\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i}),$ and $\mathcal{B}$ denotes an on-policy trajectory batch, and $\hat{A}_{i,\tau}^{\pi}$ is the estimated advantage. Gradients are backpropagated not only through the policy $\tilde{\pi}_{i}$ but also through the sampling distribution $\rho(G_{\mathcal{V}_{i}}\mid s_{\mathcal{V}_{i}})$, enabling joint optimization of action selection and interaction structure.
4.2 BayesG: Variational Latent Ego-graph Inference for Policy Optimisation
Recall that the graph distribution $\rho(G_{\mathcal{V}_{i}}\mid s_{\mathcal{V}_{i}})$, introduced in Definition 3, governs which neighbors agent $i$ attends to. To infer this stochastic distribution over the fixed physical topology, we model it via a binary mask $Z_{i}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$ applied over the physical neighborhood graph $G^{\text{env}}_{\mathcal{V}_{i}}$, where each entry $z_{ij}=1$ indicates that agent $i$ communicates with neighbor $j$. This yields the effective subgraph:
$G^{*}_{i}=Z_{i}\odot G^{\mathrm{env}}_{\mathcal{V}_{i}}.$ (6)
We formalise the learning of the latent edge mask $Z_{i}$ as a Bayesian inference problem. This is a natural choice as Bayes’ theorem offers a principled way to infer a latent distribution conditioned on the directly observable physical graph $G^{\text{env}}_{\mathcal{V}_{i}}$ and agent-local data $D_{i}$. Treating edge masks as latent variables enables agents to capture uncertainty over coordination structures—particularly valuable when noisy or non-stationary dynamics make deterministic graph pruning unreliable.
Let $D_{i}$ denote agent-local data collected during training, such as neighbor states, policy outputs, and trajectory embeddings. The posterior distribution over edge masks is given by Bayes’ theorem:
$p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})\propto p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})\cdot p(Z_{i}),$ (7)
where $p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})$ is the likelihood, measuring how well the masked graph explains observed behavior, and $p(Z_{i})$ is a prior (e.g., Bernoulli with bias).
To enable tractable optimization, we approximate the posterior with a variational distribution:
$q(Z_{i};\phi_{i})=\prod_{j\in\mathcal{N}_{i}}\text{Bern}(z_{ij};\sigma(\phi_{ij})),$ (8)
where $\phi_{i}$ are learnable logits and $\sigma(\cdot)$ is the sigmoid function. We apply the Gumbel-Softmax reparameterization[35, 36] to allow gradient-based learning through discrete edge sampling.
The variational parameters $\phi_{i}$ are learned by minimizing the Kullback–Leibler (KL) divergence between the approximate posterior $q(Z_{i};\phi_{i})$ and the true Bayesian posterior $p(Z_{i}\mid G^{\mathrm{env}}_{\mathcal{V}_{i}},D_{i})$:
$\mathrm{KL}[q(Z_{i};\phi_{i})\|p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})]=\mathbb{E}_{q(Z_{i})}\left[\log q(Z_{i})-\log p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})\right].$ (9)
This approximation enables each agent to learn a structured distribution over interaction graphs that reflects the most informative subgraphs for policy optimization. Following the standard variational inference framework, minimizing this KL divergence is equivalent (up to an additive constant) to maximizing the evidence lower bound (ELBO), or equivalently, minimizing the negative ELBO:
$\mathcal{L}_{\mathrm{ELBO}}$ $=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})+\log p(Z_{i})-\log q(Z_{i};\phi_{i})\right]+\text{const}$ (10)
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}+\log p(Z_{i})-\log q(Z_{i};\phi_{i})\right]+\text{const},$
where the likelihood term $\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})$ is instantiated by the graph-conditioned actor loss $\mathcal{L}_{\theta,\varphi}$ defined in Definition 4, which provides task feedback to guide the learning of latent structures. (See Appendix A.3 for details.)
Connection to maximum entropy RL. Our formulation parallels the probabilistic interpretation of RL in Soft Actor-Critic (SAC) [37], which casts policy optimization as entropy-regularized inference. In SAC, policy entropy $\mathcal{H}(\pi(\cdot|s))$ regularizes action selection; in our framework, the term $\log q(Z_{i};\phi_{i})$ in Equation (10) acts as mask entropy $\mathcal{H}(q(Z_{i};\phi_{i}))$, regularizing the variational distribution over communication structures. This prevents premature collapse to deterministic graphs and enables uncertainty-aware neighbor selection. The dual entropy regularization—over both actions (embedded in $\mathcal{L}_{\theta,\varphi}$) and edges (the $\log q$ term)—is key to our method’s robustness in dynamic environments where action diversity and adaptive communication are both critical.
Expanding the ELBO in Equation (10), we obtain:
Definition 5 (BayesG-ELBO Objective).
The BayesG objective integrates policy learning and graph inference in a unified variational framework. For agent $i$, the ELBO is:
$\mathcal{L}_{\mathrm{ELBO}}=\mathbb{E}_{q(Z_{i};\phi_{i})}\!\left[-\mathcal{L}_{\theta,\varphi}\right]-\sum_{j\in\mathcal{N}_{i}}\mathrm{KL}\!\left(q(Z_{ij};\phi_{ij})\|p(Z_{ij})\right),$ (11)
where $\mathcal{L}_{\theta,\varphi}$ denotes the graph-conditioned policy loss defined in Definition 4, $q(Z_{ij};\phi_{ij})=\mathrm{Bernoulli}(\sigma(\phi_{ij}))$ is the variational edge distribution, and $p(Z_{ij})=\mathrm{Bernoulli}(\lambda)$ is the Bernoulli prior with retention bias $\lambda$. This objective jointly optimizes the policy and the structure of the latent subgraph via variational inference.
We complement the actor-based ELBO with a standard A2C critic loss:
$\mathcal{L}_{\text{Critic}}=\frac{1}{|\mathcal{B}|}\sum_{\tau\in\mathcal{B}}\left(V_{\omega_{i}}(\tilde{s}_{i,\tau},u_{\mathcal{N}_{i},\tau})-\hat{R}_{i,\tau}^{\pi}\right)^{2},$ (12)
where $\tilde{s}_{i,\tau}$ is the graph-conditioned input for the critic, and $\hat{R}_{i,\tau}^{\pi}$ is the estimated return. The final training objective combines both components across agents: $\mathcal{L}_{\text{total}}=\sum_{i\in\mathcal{V}}\left(-\mathcal{L}_{\mathrm{ELBO}}+\mathcal{L}_{\text{Critic}}\right).$ This formulation enables decentralized agents to learn both communication structure and policy behavior from local interactions, while adhering to the physical structure of the environment. The ego-graph terminology reflects that each agent independently infers its own local communication neighborhood. Full derivations are provided in Appendix A.3.
5 Experiments
5.1 Experimental Setup
We evaluate BayesG on five benchmark scenarios for adaptive traffic signal control (ATSC), implemented using the SUMO microscopic traffic simulator [38]. Each environment simulates peak-hour traffic, with one MDP step corresponding to a fixed control interval. Agents observe traffic conditions via induction-loop detectors (ILDs), including vehicle density, queue lengths, and waiting times on incoming lanes, and control local traffic signals. The reward is the negative number of halted vehicles, normalized by a fixed scale. These environments naturally satisfy Definition 1: state transitions depend on local and neighbor actions, rewards are localized per agent, and neighborhood structure is fixed by the road network. Detailed MDP mappings are in Appendix B.  [^1] Code and data are available at https://github.com/Wei9711/BayesG.
ATSC_Grid and Monaco. These two medium-scale scenarios are widely used in prior ATSC benchmarks (e.g., NeurComm [7]). ATSC_Grid is a synthetic $5\times 5$ network with regular connectivity, while Monaco replicates a real-world 28-intersection layout. Both use a 5-second control interval and run for 720 steps per episode, totaling 3600 seconds of simulated time. States include lane-level traffic conditions and neighbor observations. Actions correspond to local phase switches, and a 2-second yellow phase is enforced for safety.
NewYork. To assess scalability, we introduce three large-scale scenarios—NewYork33, NewYork51, and NewYork167—comprising 33, 51, and 167 signalized intersections, respectively. These networks are derived from real-world Manhattan layouts (see Appendix C for more details). We use a 20-second control interval (increased to 40 seconds for NewYork167) and simulate 500 steps per episode. Intersections feature heterogeneous phase designs and ILDs configurations. States include normalized lane-level metrics and neighborhood-aware features. These environments pose significant challenges for coordination and generalization in decentralized MARL.
5.1.1 Baselines
We compare BayesG against six representative multi-agent actor–critic baselines, all implemented using a unified A2C backbone with consistent neighbor access and decentralized execution:
IA2C [13]: Independent actor–critic training where each agent optimizes its policy based on local observations and a critic that observes neighboring actions. No inter-agent communication is used.
ConseNet [16]: A consensus-based method where critics synchronize neighbours’ parameters via local averaging to promote stability and coordination, without exchanging message content.
FPrint [39]: Mitigates non-stationarity by attaching policy fingerprints to local transitions, enabling agents to condition their updates on the behavior of neighbors over time.
LToS [19]: Introduces a hierarchical reward-sharing scheme where agents learn to shape local rewards via neighborhood-level value estimates. No explicit communication is involved.
CommNet [40]: A communication-based architecture where each agent receives the average of all encoded neighbor messages, limiting expressiveness but allowing simple message integration.
NeurComm [7]: A more flexible communication protocol that encodes and aggregates messages from neighbors without averaging, supporting richer interaction modeling. It generalizes earlier approaches such as CommNet and DIAL [41].
We categorize IA2C, ConseNet, FPrint, and LToS as non-communicative baselines, as they do not involve direct message exchange. In contrast, CommNet, NeurComm, and our method BayesG are communication-based.
5.1.2 Implementation Details
For all environments, we use a fixed control interval of 5 seconds. Policy and critic networks share similar architectures across all baselines. Each experiment is averaged over 5 random seeds.
[…]
CommNet. Each agent encodes its own and neighboring observations with an MLP, and aggregates neighbors’ trajectory features via mean pooling followed by another MLP. The final input is the sum of these two components: $\tilde{s}_{i}=\mathrm{MLP}_{\text{state}}([s_{i},s_{\mathcal{N}_{i}}])+\mathrm{MLP}_{\text{traj}}(\text{mean}(h_{\mathcal{N}_{i}})).$
NeurComm. NeurComm encodes each of the three information types separately and concatenates their embeddings before passing the result into an LSTM for temporal reasoning: $\tilde{s}_{i}=\mathrm{MLP}_{\text{state}}([s_{i},s_{\mathcal{N}_{i}}])\,\|\,\mathrm{MLP}_{\text{policy}}(\pi_{\mathcal{N}_{i}})\,\|\,\mathrm{MLP}_{\text{traj}}(h_{\mathcal{N}_{i}}).$
BayesG (Ours). BayesG replaces fixed communication with a learned latent subgraph. Let $A_{i}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$ denote the local physical graph and $Z_{i}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$ a sampled binary mask; the effective graph is defined as: $A_{i}^{*}=Z_{i}\odot A_{i}.$ The input is then computed via masked message passing over $A_{i}^{*}$ using GNNs for each input channel:
$\tilde{s}_{i}=\mathrm{GNN}_{\text{obs}}(S_{\mathcal{V}_{i}},A_{i}^{*})\,\|\,\mathrm{GNN}_{\text{policy}}(\Pi_{\mathcal{V}_{i}},A_{i}^{*})\,\|\,\mathrm{GNN}_{\text{traj}}(H_{\mathcal{V}_{i}},A_{i}^{*}).$ (13)
Here, $S_{\mathcal{V}_{i}}$, $\Pi_{\mathcal{V}_{i}}$, and $H_{\mathcal{V}_{i}}$ represent the observation, policy, and trajectory features of agent $i$’s neighborhood. While any GNN can be employed, we use graph convolutional networks (GCNs) [42] for efficiency in our implementation. This formulation allows each agent to learn sparse and context-aware communication structures tailored to local dynamics. (See pseudocode in Appendix D).
Figure 2: Training reward curves of BayesG and baselines across five ATSC environments. BayesG consistently achieves higher returns and faster convergence, particularly in large-scale settings, demonstrating the benefit of learning task-adaptive communication graphs.
5.2 Performance Comparison
Training Performance. Figure 2 presents the training curves of BayesG and baseline methods across five ATSC environments. Each curve shows the smoothed average episode return (defined as the negative of total queue length) over 5 random seeds. Several key trends emerge: (1) Consistent superiority of BayesG: Across all environments, BayesG steadily outperforms both non-communicative baselines (IA2C, ConseNet, FPrint, LToS) and explicit communication protocols (CommNet, NeurComm). This highlights the advantage of learning task-adaptive latent subgraphs that promote efficient and stable coordination. (2) Faster convergence: BayesG exhibits faster convergence in early training, particularly on ATSC_Grid and Monaco. This indicates that selective interaction via learned masks accelerates policy learning by reducing noisy or redundant communication. (3) Scalability in large networks: On the large-scale NewYork33, NewYork51, and NewYork167 scenarios, BayesG achieves higher asymptotic returns and more stable learning curves, while baselines often suffer from plateauing or instability. This validates BayesG’s ability to scale to high-dimensional decentralized settings by suppressing uninformative edges. These results collectively demonstrate that BayesG not only enhances learning efficiency but also improves final policy quality by discovering sparse, context-aware coordination patterns.
To further understand the optimization dynamics of BayesG, we report the evolution of key training loss components in Appendix E for ATSC_Grid, Monaco, and NewYork33. Specifically, we track the policy loss, value (critic) loss, ELBO loss, prior loss, mask entropy regularization, and the total objective. These loss curves offer deeper insight into how BayesG balances policy optimization with variational graph inference, and demonstrate the stability of the joint training process across different network scales.
Figure 3: Visualization of traffic congestion on the Grid map at 3500 simulation seconds. Road thickness and color represent vehicle density (thicker and redder indicates more congestion). BayesG results in significantly lighter and sparser traffic flows.Additional qualitative comparisons across different simulation times are provided in Figure 10 (Appendix F).
Qualitative Comparison on Grid Environment. Figure 3 presents a visual comparison of traffic congestion for IA2C, CommNet, NeurComm, and BayesG on the ATSC_Grid map at 3,500 simulation seconds. Road segment thickness and color indicate vehicle density, with thicker and redder lines denoting heavier congestion. IA2C and CommNet exhibit severe congestion across many intersections, while NeurComm reduces some bottlenecks through explicit communication. BayesG, in contrast, achieves the smoothest traffic flow, with thinner and less congested segments. These results highlight BayesG’s ability to adaptively focus on critical interactions via its learned latent graph, enabling more effective and scalable coordination in dense traffic scenarios. Additional qualitative comparisons across different simulation times are provided in Appendix F.
Figure 4: Case study on the ATSC_Grid map at time step 1400. (a) The latent interaction graph inferred by BayesG. Each agent samples a probabilistic binary mask over its ego-graph and the global latent graph is formed by aggregating these per-agent ego-graph masks. Edge thickness reflects the inferred likelihood of communication between intersections. (b) The vehicle density snapshot from the SUMO simulation. Thicker red/orange lines denote higher vehicle accumulation on road segments, indicating local congestion.
5.3 Case Study: Interpreting Learned Interaction Structures
To further understand the behavior of BayesG, we conduct a case study on the ATSC_Grid map at time step 1,400. Figure 4 illustrates: (a) Left: The latent interaction matrix, where each row corresponds to an agent’s inferred ego-graph—i.e., a probabilistic mask over its physical neighborhood. The aggregated matrix represents the likelihood of interaction between all agent pairs.(a) Right: The physical road network layout with edges colored and weighted according to the same communication probabilities, providing an interpretable spatial view. (b): The vehicle density snapshot from SUMO simulation at the same time step. Thicker and redder road segments indicate higher congestion levels.
This visualization shows that BayesG learns to prioritize communication from low-congestion intersections toward more congested regions. For example, when congestion builds up in a particular area (e.g., top-right and bottom right clusters), left-stream intersections increase their coordination weights, effectively regulating inflow and providing more green phases to help alleviate downstream pressure. Unlike approaches that treat all neighbors equally, BayesG captures directional, task-driven cooperation: it encourages agents outside a congested zone to adjust proactively, thereby relieving bottlenecks before they worsen. This reflects real-world traffic control principles, where strategic upstream coordination plays a critical role in minimizing congestion spread. These results demonstrate BayesG’s ability to infer decentralized, context-aware communication structures that adapt dynamically to evolving traffic conditions. (See more cases in Appendix G)
Figure 5: Ablation study on the Monaco and NewYork51 environments. Left: Performance comparison of different graph masking strategies. Right: Effect of different feature types used to generate the variational mask.
5.4 Ablation Studies
To better understand the impact of BayesG’s components, we conduct two sets of ablation studies on the Monaco and NewYork51 environments, focusing on (i) how the graph mask is generated, and (ii) what input features are used to learn the mask. Results are summarized in Figure 5.
Graph Masking Strategies. We compare three strategies for subgraph construction: (1) No Masking: Uses the raw environment topology $A_{i}$ without any pruning. (2) Random Masking: Applies a randomly sampled binary mask $Z_{i}$, simulating naive edge dropout. (3) Learned Mask (BayesG): Applies the full variational inference pipeline to infer $Z_{i}$, resulting in adaptive and task-specific subgraphs. As shown in Figure 5 (left), random masking severely degrades performance, while the learned mask significantly outperforms both baselines by suppressing irrelevant links and preserving only informative neighbor connections.
Mask Input Features. We further examine how different information types affect the learned mask. Specifically, we use:(1) State: Mask is generated based on neighbor observations $s_{\mathcal{V}_{i}}$. (2) Trajectory: Uses LSTM hidden states $h_{\mathcal{V}_{i}}$, capturing temporal behavior. (3) Policy: Based on neighbor action distributions $\pi_{\mathcal{N}_{i}}$. Results in Figure 5 (right) indicate that trajectory-based features lead to stronger performance that state-only, as the incorporate richer temporal context. Policy-based masks also provide meaningful signals, reflecting agents’ behavioral intentions. The best results are achieved when the learned mask uses a combination of these features, as implemented in BayesG.
6 Discussion
6.1 Comparison with Sampling-based Networked MARL
Recent work has explored sampling strategies for scalability in networked MARL, but differs fundamentally in what is sampled, how distributions are determined, and where adaptivity occurs.
Sampling targets and distributions. Qu et al. [26] establish convergence results for networked systems with local dependencies on fixed graphs. Lin et al. [23] sample active links from a predefined distribution $\mathcal{D}$ for $\mu$-decay analysis. Anand and Qu [27] use uniform agent subsampling ($k$ out of $n$) for a global controller. In contrast, our method learns context-dependent, per-edge distributions $q(Z_{i};\phi_{i})$ optimized jointly with the policy via an ELBO.
Locus of adaptivity. Lin et al. [23] achieve dynamics-level stochasticity through time-varying link sets. Anand and Qu [27] vary agent subsets for a global controller. Our approach operates at the policy level: each agent samples a task-adaptive 1-hop subgraph for local decision-making.
First-hop inclusion probabilities. Lin et al. [23] use the marginal of $\mathcal{D}$ (predefined, not state-adaptive). Anand and Qu [27] use uniform inclusion ($k/n$). Our method learns edge-specific probabilities $\Pr[z_{ij}=1]=\sigma(\phi_{ij}(\cdot))$ that are state and trajectory-dependent.
These approaches provide valuable convergence guarantees under structural assumptions. Our contribution is complementary: learning which neighbors to attend to, optimized end-to-end with the policy. A detailed comparison is in Table 2 (Appendix H).
6.2 Distinction from Dec-POMDP-based Coordination Graphs
Our method addresses the Spatiotemporal-MDP setting (Definition 1), where agent transitions $p_{i}(s_{i}^{\prime}|s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}})$ and rewards $R(s_{\mathcal{V}_{i}},u_{\mathcal{V}_{i}})$ are localized over physically connected neighbors determined by $G_{\text{env}}$. This reflects real-world systems (traffic networks, power grids, sensor fields) where agents only observe and affect local neighborhoods through fully decentralized learning.
In contrast, Dec-POMDP-based methods such as CASEC [15], DCG [33], DICG [29], SOP-CG [34], and GACG [22] assume global state $s$, global reward $R(s,\mathbf{u})$, joint transitions $p(s^{\prime}|s,\mathbf{u})$, and centralized training with decentralized execution (CTDE). They allow unrestricted graph rewiring unconstrained by physical topology, making them incompatible with physically grounded networked systems. Our baselines—CommNet [40], NeurComm [7], and LToS [23]—communicate only with physical neighbors, respect fixed topology, and perform fully decentralized learning. While Dec-POMDP coordination graphs are important contributions, we exclude them due to incompatibility with Spatiotemporal-MDP’s decentralized, physically constrained nature.
7 Conclusion
We presented BayesG, a variational framework for learning latent interaction graphs in networked multi-agent reinforcement learning. By operating over ego-graphs—localized subgraphs centered on each agent—BayesG enables decentralized agents to infer task-adaptive communication structures using local observations and Bayesian inference. Our approach integrates graph inference with actor–critic training via an ELBO objective, jointly optimizing both the interaction topology and policy. Experiments on five adaptive traffic signal control benchmarks demonstrate that BayesG consistently outperforms strong baselines, achieving superior coordination and reduced congestion. Case studies and ablation results further highlight the efficiency and interpretability of the learned graphs. BayesG offers a principled approach to structure-aware cooperation in decentralized settings and opens up new opportunities for scalable learning in networked systems.
Acknowledgements
This work is supported by the Australian Research Council under Australian Laureate Fellowships FL190100149.
Appendix
Appendix A Detailed Derivation
A.1 Derivation of Graph-based Policy
We expand on Definition 3, which defines each agent's policy as a two-stage process: (i) sampling a binary subgraph $G_{\mathcal{V}_{i}}$ from a variational distribution $\rho$, and (ii) conditioning the action policy on the graph-filtered encoding $\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}})$.
In conventional MARL, each agent conditions its policy on a fixed observation $\tilde{s}_{i}=f(s_{\mathcal{V}_{i}})$, where $\mathcal{V}_{i}$ is the agent’s closed neighborhood. In our formulation, the observation additionally depends on a latent subgraph $G_{\mathcal{V}_{i}}$ drawn from a learned distribution:
$G_{\mathcal{V}_{i}}\sim\rho(G\mid s_{\mathcal{V}_{i}};\varphi_{i}),$ (14)
where $\rho$ is parameterized by variational parameters $\varphi_{i}$. To ensure feasibility, we constrain sampled subgraphs to lie within the physical topology:
$G_{\mathcal{V}_{i}}\preceq G^{\text{env}}_{\mathcal{V}_{i}},$ (15)
i.e., agents can only sample from edges permitted by the environment graph.
The graph-conditioned encoder $\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}})$ is implemented using GNNs over the masked subgraph. The resulting policy is defined as:
$\pi_{i}(u_{i}\mid s_{\mathcal{V}_{i}})=\mathbb{E}_{G_{\mathcal{V}_{i}}\sim\rho}\left[\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}});\theta_{i})\right],$ (16)
where $\tilde{\pi}_{i}$ is the action policy given graph-filtered input.
This formulation enables each agent to dynamically adapt its local observation space via sampled interaction subgraphs, supporting sparse and context-aware reasoning under topological constraints.
A.2 Derivation of Graph-based A2C Objective
We derive the actor loss introduced in Definition 4, which integrates latent subgraph sampling into A2C learning.
Recall the standard A2C actor loss:
$\mathcal{L}^{\text{A2C}}(\theta_{i})=-\log\pi_{\theta_{i}}(u_{i}\mid\tilde{s}_{i})\cdot\hat{A}_{i}^{\pi},$ (17)
where $\tilde{s}_{i}$ is the observation, $u_{i}$ the action, and $\hat{A}_{i}^{\pi}$ the estimated advantage.
Under the graph-based policy framework, the policy depends on a sampled subgraph $G_{\mathcal{V}_{i}}$:
$\pi_{i}(u_{i}\mid s_{\mathcal{V}_{i}})=\mathbb{E}_{G_{\mathcal{V}_{i}}\sim\rho(\cdot\mid s_{\mathcal{V}_{i}};\varphi_{i})}\left[\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}});\theta_{i})\right].$ (18)
The expected actor loss becomes:
$\mathcal{L}_{\theta,\varphi}=\mathbb{E}_{G_{\mathcal{V}_{i}}\sim\rho}\left[-\log\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}}))\cdot\hat{A}_{i}^{\pi}\right].$ (19)
We add entropy regularization to encourage exploration:
$\mathcal{H}(\tilde{\pi}_{i}(\cdot\mid\tilde{f}_{i}))=-\sum_{u_{i}\in\mathcal{U}^{i}}\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i})\log\tilde{\pi}_{i}(u_{i}\mid\tilde{f}_{i}).$ (20)
Over a batch $\mathcal{B}$, the full objective is:
$\mathcal{L}_{\theta,\varphi}=\frac{1}{|\mathcal{B}|}\sum_{\tau\in\mathcal{B}}\mathbb{E}_{G_{\mathcal{V}_{i}}\sim\rho}\left[-\log\tilde{\pi}_{i}(a_{i,\tau}\mid\tilde{f}_{i}(s_{\mathcal{V}_{i}},G_{\mathcal{V}_{i}}))\cdot\hat{A}_{i,\tau}^{\pi}+\beta\cdot\mathcal{H}(\tilde{\pi}_{i}(\cdot\mid\tilde{f}_{i}))\right].$ (21)
This defines the graph-based actor loss used in our main training objective.
A.3 Derivation of Variational Objective and ELBO for BayesG
We present the detailed derivation of the ELBO objective used in BayesG, beginning with Bayes’ theorem and ending with a tractable training objective for policy and graph optimization.
Bayesian Inference over Latent Graphs
In networked MARL, the environment graph $G^{\text{env}}_{\mathcal{V}_{i}}$ represents the physical connectivity between agents (e.g., traffic lights connected via roads), which imposes hard topological constraints on interaction. We model the stochastic interaction graph for agent $i$ as a binary mask $Z_{i}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$, sampled from a learned distribution over the physical neighborhood graph $G^{\text{env}}_{\mathcal{V}_{i}}$. The resulting posterior distribution is:
$p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})=\frac{p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})\cdot p(Z_{i})}{p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})},$ (22)
where:
  • $D_{i}$ denotes the agent-specific data (e.g., neighbor states, trajectories, and policies),
  • $p(Z_{i})$ is a prior over edge masks,
  • $p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})$ is the likelihood of the data under the masked graph,
  • $p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})$ is the marginal likelihood.
Variational Approximation and KL Divergence
Since the exact posterior in Eq. (22) is intractable, we introduce a variational distribution $q(Z_{i};\phi_{i})$ to approximate it. The variational parameters $\phi_{i}$ are learned by minimizing the Kullback–Leibler (KL) divergence between the approximate posterior and the true posterior:
$\mathrm{KL}[q(Z_{i};\phi_{i})\|p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})]=\mathbb{E}_{q(Z_{i})}\left[\log q(Z_{i})-\log p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})\right].$ (23)
Applying Bayes’ rule to the log posterior, we rewrite:
$\log p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})=\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})+\log p(Z_{i})-\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}}).$ (24)
Substituting Eq. (24) into Eq. (23) yields:
$\mathrm{KL}$ $=\mathbb{E}_{q(Z_{i})}\left[\log q(Z_{i})-\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})-\log p(Z_{i})+\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})\right]$ (25)
$=\mathbb{E}_{q(Z_{i})}\left[\log q(Z_{i})-\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})-\log p(Z_{i})\right]+\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})$ (26)
$=-\mathbb{E}_{q(Z_{i})}\left[\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})+\log p(Z_{i})-\log q(Z_{i})\right]+\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})$ (27)
$=-\mathbb{E}_{q(Z_{i})}\left[\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})+\log p(Z_{i})-\log q(Z_{i})\right]+\text{const},$ (28)
where we have used the fact that $\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})$ does not depend on $Z_{i}$ or $\phi_{i}$. Therefore, minimizing the KL divergence is equivalent (up to an additive constant) to maximizing the evidence lower bound (ELBO):
$\mathcal{L}_{\mathrm{ELBO}}=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[\log p(D_{i}\mid Z_{i})+\log p(Z_{i})-\log q(Z_{i};\phi_{i})\right]+\text{const}.$ (29)
Term-by-Term Breakdown
1. Likelihood Term: $\log p(D_{i}\mid Z_{i})$
This is modeled using the graph-conditioned policy loss, i.e.,
$\log p(D_{i}\mid Z_{i})\approx-\mathcal{L}_{\theta,\varphi},$ (30)
where $\mathcal{L}_{\theta,\varphi}$ is the actor loss under sampled subgraph $Z_{i}\odot G^{\text{env}}_{\mathcal{V}_{i}}$.
2. Prior Term: $\log p(Z_{i})$
We define the prior as an element-wise Bernoulli with retention bias $\lambda$:
$p(Z_{i})=\prod_{j\in\mathcal{N}_{i}}\lambda^{z_{ij}}(1-\lambda)^{1-z_{ij}}.$ (31)
Then:
$\log p(Z_{i})=\sum_{j\in\mathcal{N}_{i}}z_{ij}\log\lambda+(1-z_{ij})\log(1-\lambda).$ (32)
Taking expectation under $q(Z_{i})$:
$\mathbb{E}_{q}[\log p(Z_{i})]=\sum_{j\in\mathcal{N}_{i}}\left[\sigma(\phi_{ij})\log\lambda+(1-\sigma(\phi_{ij}))\log(1-\lambda)\right].$ (33)
3. Entropy Term: $\log q(Z_{i};\phi_{i})$
Since $q(Z_{i})$ is a factorized Bernoulli:
$H(q(Z_{ij}))=-\sigma(\phi_{ij})\log\sigma(\phi_{ij})-(1-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij})).$ (34)
Then:
$\mathbb{E}_{q}[\log q(Z_{i})]=-\sum_{j\in\mathcal{N}_{i}}H(q(Z_{ij})).$ (35)
Final Objective
Combining all terms:
$\mathcal{L}_{\mathrm{ELBO}}$ $=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]+\sum_{j\in\mathcal{N}_{i}}\left[\sigma(\phi_{ij})\log\lambda+(1-\sigma(\phi_{ij}))\log(1-\lambda)\right]+\sum_{j\in\mathcal{N}_{i}}H(q(Z_{ij}))$ (36)
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]+\sum_{j\in\mathcal{N}_{i}}\left[\sigma(\phi_{ij})\log\lambda+(1-\sigma(\phi_{ij}))\log(1-\lambda)\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{j\in\mathcal{N}_{i}}\left[-\sigma(\phi_{ij})\log\sigma(\phi_{ij})-(1-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij}))\right]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\Bigg[-\mathcal{L}_{\theta,\varphi}+\sum_{j\in\mathcal{N}_{i}}\Big(\sigma(\phi_{ij})\log\lambda+(1-\sigma(\phi_{ij}))\log(1-\lambda)$
$\qquad\qquad\qquad\qquad\qquad\qquad-\sigma(\phi_{ij})\log\sigma(\phi_{ij})-(1-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij}))\Big)\Bigg]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}+\sum_{j\in\mathcal{N}_{i}}\left(\sigma(\phi_{ij})\log\frac{\lambda}{\sigma(\phi_{ij})}+(1-\sigma(\phi_{ij}))\log\frac{1-\lambda}{1-\sigma(\phi_{ij})}\right)\right]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]-\sum_{j\in\mathcal{N}_{i}}\mathrm{KL}\left(q(Z_{ij};\phi_{ij})\|p(Z_{ij})\right),$
This is the BayesG training objective as formalized in Definition 5.
Figure 6: Visualization of the NewYork33, NewYork51, and NewYork167 environments. Top: SUMO network with signalized intersections. Bottom: extracted graph structure used in networked MARL, including traffic light nodes and their physical neighbors.
Appendix B ATSC Environment: MDP Component Mapping and Spatiotemporal-MDP Justification
We provide detailed mappings from the adaptive traffic signal control (ATSC) domain to the Spatiotemporal-MDP framework (Definition 1).
B.1 MDP Components
States ($s_{i}$). Each agent $i$ (a signalized intersection) observes its local traffic state via induction loop detectors (ILDs) placed on incoming lanes. The state includes:
  • Vehicle density: Number of vehicles per lane
  • Queue length: Number of stopped vehicles per lane
  • Waiting time: Average waiting time of vehicles per lane
  • Current phase: Active traffic signal phase (e.g., north-south green)
  • Phase duration: Time elapsed in current phase
For neighborhood-aware coordination, agents also receive aggregated statistics from immediate neighbors $\mathcal{N}_{i}$ (e.g., neighbor queue lengths, active phases).
Actions ($u_{i}$). Each agent selects a traffic signal phase from a discrete action space. Actions correspond to:
  • Phase switching: Transition to a different signal phase (e.g., from north-south green to east-west green)
  • Phase holding: Maintain the current phase for another control interval
A mandatory yellow phase (2 seconds) is enforced between conflicting green phases for safety.
Transition Probabilities ($p_{i}(s_{i}^{\prime}|s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}})$). Transitions are not analytically available but are governed by the SUMO microscopic traffic simulator [38]. The key property is locality: agent $i$’s next state $s_{i}^{\prime}$ depends primarily on:
  • Its own action $u_{i}$ (phase decision)
  • Local state $s_{i}$ (current congestion)
  • Neighbor actions $u_{\mathcal{N}_{i}}$ (upstream/downstream traffic release)
This satisfies the Spatiotemporal-MDP assumption that transitions are localized over the physical neighborhood $\mathcal{V}_{i}$.
Rewards ($R(s_{\mathcal{V}_{i}},u_{\mathcal{V}_{i}})$). The reward for agent $i$ is:
$r_{i}=-\frac{1}{C}\sum_{\ell\in\text{lanes}_{i}}\text{num\_halted}(\ell),$ (37)
where $\text{num\_halted}(\ell)$ is the number of stopped vehicles on lane $\ell$, and $C$ is a normalization constant. This encourages local queue minimization. The reward is localized to agent $i$’s intersection, consistent with the Spatiotemporal-MDP framework.
B.2 Why Spatiotemporal-MDP Fits ATSC
The ATSC domain is a canonical example of Spatiotemporal-MDP due to:
1. Localized dynamics. Traffic flow is governed by physical proximity: upstream intersections release vehicles that propagate to downstream intersections. Each agent’s state evolution depends on its immediate neighbors’ actions, not the global joint action of all agents.
2. Fixed physical topology. The road network structure is fixed and sparse, with agents (intersections) only interacting with directly connected neighbors via shared road segments.
3. Decentralized execution requirement. In real-world deployments, traffic signals operate independently with limited communication bandwidth. Centralized control is impractical due to:
  • Scalability: City-scale networks have hundreds of intersections; centralized joint action spaces grow exponentially
  • Communication constraints: Real-time global state aggregation is infeasible under latency and bandwidth limits
  • Robustness: Centralized systems are vulnerable to single points of failure
4. Local observability. Each intersection has sensors only for its incoming lanes, consistent with the partial observability assumption in Spatiotemporal-MDP.
These properties make ATSC fundamentally different from cooperative benchmarks (e.g., StarCraft) that assume global rewards, unrestricted communication, and arbitrary coordination graphs. Our method exploits this structure by learning sparse, physically grounded communication masks, enabling scalable and deployable traffic control.
Appendix C NewYork Scenario Visualization and Graph Statistics
To support the evaluation of BayesG on large-scale environments, we include visualizations and statistics for the three real-world maps: NewYork33, NewYork51, and NewYork167. These maps are derived from SUMO simulations based on real intersections in Manhattan, and are visualized in Figure 6.
Each row in Figure 6 shows:
  • Top: The SUMO map used for microscopic traffic simulation, where green nodes represent traffic light-controlled intersections.
  • Bottom: The corresponding environment graph used for MARL training. Blue circles represent agents (traffic lights), and red edges indicate neighbor connections used for policy input.
Note: The graph used in our MARL environment is a subgraph of the physical road network. Specifically, we include only intersections that are signal-controlled (i.e., managed by traffic lights), and define edges based on direct traffic flow connections between these nodes. Road segments that connect to unsignalized intersections or that skip over intermediate traffic lights are excluded from the environment graph. This design reflects the decentralized setting in which agents can only communicate and coordinate with neighboring controlled intersections.
Scenario # Nodes # Edges Avg. Degree Max Degree
NewYork33 33 56 1.70 3
NewYork51 51 125 2.45 4
NewYork167 167 384 2.30 4
Table 1: Graph statistics for the NewYork traffic signal control environments. Nodes represent signalized intersections; edges represent physical connectivity between controlled agents.
Appendix D Algorithms
We provide pseudocode for BayesG training and decentralized execution. Algorithm 1 outlines asynchronous multi-agent training with latent graph inference, following the procedures in Sections 3 and 4. The training loop includes: (i) edge sampling and message propagation, (ii) policy and trajectory updates, (iii) value estimation and environment simulation, and (iv) gradient-based updates of actor, critic, and variational graph parameters.
Algorithm 2 describes decentralized execution: each agent samples a subgraph, encodes its neighborhood, and selects actions in real time without global information.
Algorithm 1 BayesG: Multi-agent A2C Training with Variational Graph Inference
1:Parameter: $\alpha,\beta,\gamma,T,|\mathcal{B}|,\eta_{\theta},\eta_{\omega},\eta_{\phi}$
2:Result: $\{\theta_{i},\omega_{i},\phi_{i}\}_{i\in\mathcal{V}}$
3:Initialize $s_{0}$, $\pi_{-1}$, $h_{-1}$, $t\leftarrow 0$, $k\leftarrow 0$, $\mathcal{B}\leftarrow\emptyset$
4:repeat
5:  for $i\in\mathcal{V}$ do
6:   Sample $Z_{i}\sim q(Z_{i};\phi_{i})$, set $A_{i}^{*}\leftarrow Z_{i}\odot G^{\mathrm{env}}_{\mathcal{V}_{i}}$
7:   Encode $\tilde{s}_{i,t}\leftarrow\tilde{f}_{i}(s_{\mathcal{V}_{i}},A_{i}^{*})$
8:   Update $h_{i,t}\leftarrow g_{\nu_{i}}(h_{i,t-1},\tilde{s}_{i,t})$
9:   Sample $a_{i,t}\sim\pi_{\theta_{i}}(\cdot\mid h_{i,t})$
10:   Compute $v_{i,t}\leftarrow V_{\omega_{i}}(h_{i,t},u_{\mathcal{N}_{i},t})$
11:   Execute $a_{i,t}$
12:  end for
13:  Simulate $\{s_{i,t+1},r_{i,t}\}_{i\in\mathcal{V}}$
14:  $\mathcal{B}\leftarrow\mathcal{B}\cup\{(s_{i,t},\pi_{i,t-1},a_{i,t},r_{i,t},v_{i,t})\}_{i\in\mathcal{V}}$
15:  $t\leftarrow t+1$, $k\leftarrow k+1$
16:  if $t=T$ then
17:   Reset $s_{0}$, $\pi_{-1}$, $h_{-1}$, $t\leftarrow 0$
18:  end if
19:  if $k=|\mathcal{B}|$ then
20:   for $i\in\mathcal{V}$ do
21:     Estimate $\hat{R}_{i,\tau}^{\pi}$, $\hat{A}_{i,\tau}^{\pi}$, $\forall\tau\in\mathcal{B}$
22:     Update critic: $\omega_{i}\leftarrow\omega_{i}-\eta_{\omega}\nabla\mathcal{L}_{\text{Critic}}$
23:     Update actor & graph: $\theta_{i},\phi_{i}\leftarrow\theta_{i},\phi_{i}-\eta_{\theta}\nabla_{\theta_{i},\phi_{i}}(-\mathcal{L}_{\text{ELBO}})$
24:   end for
25:   Reset $\mathcal{B}\leftarrow\emptyset$, $k\leftarrow 0$
26:  end if
27:until Stop condition is reached
Algorithm 2 BayesG: Multi-agent Execution with Latent Graph Sampling
1:Parameter: $\{\theta_{i},\nu_{i},\phi_{i}\}_{i\in\mathcal{V}},\Delta t_{\text{comm}},\Delta t_{\text{control}}$
2:for $i\in\mathcal{V}$ do
3:  Initialize $h_{i}\leftarrow 0$, $\pi_{i}\leftarrow 0$, $\{s_{j},\pi_{j}\}_{j\in\mathcal{N}_{i}}\leftarrow 0$
4:  repeat
5:   Observe $s_{i}$
6:   Sample mask $Z_{i}\sim q(Z_{i};\phi_{i})$, compute $A_{i}^{*}\leftarrow Z_{i}\odot G^{\mathrm{env}}_{\mathcal{V}_{i}}$
7:   Send $s_{i},\pi_{i}$ to neighbors
8:   for $j\in\mathcal{N}_{i}$ do
9:     Receive and update $s_{j},\pi_{j}$ within $\Delta t_{\text{comm}}$
10:   end for
11:   Construct graph-conditioned input: $\tilde{s}_{i}\leftarrow\tilde{f}_{i}(s_{\mathcal{V}_{i}},A_{i}^{*})$
12:   Update $h_{i}\leftarrow g_{\nu_{i}}(h_{i},\tilde{s}_{i})$, compute policy $\pi_{i}\leftarrow\pi_{\theta_{i}}(\cdot\mid h_{i})$
13:   Execute action $a_{i}\sim\pi_{i}$
14:   Sleep for duration $\Delta t_{\text{control}}$
15:  until Stop condition is reached
16:end for
Figure 7: Training loss curves of BayesG on ATSC_Grid.
Figure 8: Training loss curves of BayesG on Monaco.
Figure 9: Training loss curves of BayesG on NewYork33.
Appendix E Training Loss Analysis
Figures 7, 8, and 9 illustrate the evolution of training losses for BayesG on three representative environments: ATSC_Grid, Monaco, and NewYork33. We report the component-wise losses across five random seeds.
Policy loss $\mathcal{L}_{\theta,\varphi}$.
The policy loss reflects the negative log-probability of selected actions, weighted by the estimated advantage (see Definition 4). During training, temporary increases in this loss may occur due to distributional shift in the graph-conditioned state as the variational mask distribution $q(Z_{i};\phi_{i})$ evolves. As the latent subgraphs adapt, agents may explore new action patterns that momentarily reduce alignment with advantage estimates, leading to transient spikes. However, as both the policy and graph inference converge, the loss gradually stabilizes, indicating improved policy learning under the inferred interaction structures.
ELBO loss $\mathcal{L}_{\mathrm{ELBO}}$.
The ELBO combines the actor loss with KL regularization terms (see Definition 5). On Grid and Monaco, we observe an increasing trend in the ELBO, which stems from the increasing policy confidence (lower entropy) and the corresponding rise in the KL penalty due to deviation from the Bernoulli prior. Notably, the ELBO remains smooth and stable across seeds.
On NewYork33, a different pattern emerges: the ELBO decreases in early training as the agent learns useful sparse subgraphs and aligns with the prior. Later, however, as the policy sharpens and latent graphs become more deterministic, the KL term increases, causing the ELBO to rise again. This U-shaped behavior highlights a natural trade-off between policy certainty and exploration through stochastic subgraph sampling.
Other losses.
The mask entropy term gradually decreases, indicating a transition from exploratory communication structures to more deterministic ones. The prior loss stabilizes, confirming convergence toward a learned sparsity level. The value loss steadily decreases, showing reliable value function learning.
These results confirm that BayesG effectively balances exploration and exploitation during latent graph learning and demonstrates stable convergence behavior across varying scales of networked environments.
Figure 10: Qualitative comparison at different simulation times (1000–3599s) on ATSC_Grid. Road thickness and color represent vehicle density (thicker and redder indicates more congestion). BayesG consistently learns to avoid congestion by adaptively shaping communication, in contrast to fixed or overly dense schemes used by baselines.
Appendix F Additional Visualizations on Grid Environment
To complement the qualitative analysis in Figure 3 of the main paper, we provide additional visualizations of traffic conditions and coordination patterns for IA2C, CommNet, NeurComm, and BayesG on the ATSC_Grid map. These snapshots are captured at multiple simulation times: 1000, 1500, 2000, 3000, and 3599 seconds.
Figure 10 displays the evolving traffic state and learned communication structure for each method. In these visualizations, the thickness and color intensity of each road segment reflect real-time vehicle density—thicker and redder segments indicate higher congestion levels.
Key observations:
  • IA2C struggles consistently across all time points due to the absence of inter-agent communication, resulting in persistent congestion.
  • CommNet shows slightly improved flow but lacks adaptivity, often over-communicating with irrelevant neighbors.
  • NeurComm mitigates congestion more effectively by selectively encoding neighbor messages, though its structure remains relatively fixed.
  • BayesG consistently exhibits the most adaptive behavior, dynamically adjusting its latent communication graph over time to emphasize critical neighbors and suppress redundant links, leading to the smoothest overall flow.
These time-resolved results further highlight the importance of dynamic, ego-graph-based communication in complex, real-world coordination tasks.
Figure 11: Extended case study visualizations for BayesG on ATSC_Grid. Each row shows a different simulation step and training run. (a) Left: Latent interaction matrix (ego-graph probabilities). (a) Right: Physical network view with edge thickness indicating communication strength.
Appendix G Extended Case Study: Dynamic Graph Adaptation Across Time and Runs
To further validate BayesG’s ability to infer dynamic and context-aware communication structures, we present additional case studies across two independent training runs on the ATSC_Grid environment (see Figure 11).
Each row in Figure 11 corresponds to a different simulation time step and training run: Run-1 shows the system behavior at step 600 and 1600, while Run-2 presents steps 1000 and 1400. For each time step, we visualize:
  • (a) Left: The aggregated latent interaction matrix, where each row corresponds to an agent’s inferred ego-graph—a probabilistic mask over its physical neighborhood. The overall matrix reflects the likelihood of communication between all agent pairs.
  • (a) Right: The corresponding spatial network visualization of the inferred graph, where edge width encodes communication probability. This provides an interpretable view of where information is likely to flow across the physical grid.
  • (b): The corresponding traffic density snapshot from the SUMO simulator. Redder and thicker lines denote higher congestion levels on road segments.
These results highlight how BayesG dynamically adjusts interaction structures based on evolving traffic conditions. In both runs, we observe that as local congestion increases, nearby agents increase their communication probabilities—effectively directing more coordination toward problematic areas. This behavior illustrates the emergence of adaptive, directional cooperation: upstream or adjacent intersections proactively modulate traffic signals to help alleviate pressure downstream. Such behavior aligns with human-designed traffic heuristics, where context-sensitive cooperation is essential for preventing cascading congestion.
Appendix H Detailed Comparison with Sampling-based Networked MARL
Table 2 provides a detailed side-by-side comparison of our approach with recent sampling-based methods for networked MARL [23, 27]. The comparison highlights key differences in sampling targets, distribution types, locus of adaptivity, and first-hop inclusion probabilities.
Table 2: Comparison of sampling strategies in networked MARL. Our method learns context-dependent, per-edge distributions optimized jointly with the policy, in contrast to predefined or uniform sampling approaches.
Aspect Lin et al. [23] Anand & Qu [27] BayesG (Ours)
Distribution for sampling Exogenous, fixed distribution $\mathcal{D}$ over active link sets $(L_{t}^{s},L_{t}^{r})$, sampled i.i.d.; assumes distance truncation/decay Fixed, uniform distribution over size-$k$ subsets of local agents for value estimation and action Learned, per-agent variational posterior $q(Z_{i};\phi_{i})$ over 1-hop edges; state/trajectory-dependent, trained end-to-end (ELBO)
Sampling target Global edges that are "active" for transitions ($L^{s}$) and rewards ($L^{r}$) at each step Agents (a subset of locals) for the single global agent to aggregate over (not edges) Binary edge mask on each agent’s 1-hop ego-graph (which neighbors to "listen to" for policy/communication)
Locus of dynamic topology Dynamics-level stochasticity: new active link sets each step; multi-hop influence via chains across time Agent-level subsampling: the global agent’s view changes as different locals are sampled; topology effectively star-like, not edge-level Policy/representation-level adaptivity: each agent samples a task-adaptive 1-hop subgraph each step; no multi-hop rewiring (multi-hop only via repeated local passes)
1-hop inclusion probability Inclusion of $(j\!\to\!i)$ is the marginal of $\mathcal{D}$ (fixed/model-assumed; not learned or state-adaptive) No edge-level 1-hop probability; agent-level inclusion is $k/n$ (uniform, fixed) per step for the global agent For neighbor $j$ of agent $i$, inclusion $\Pr[z_{ij}{=}1]=\sigma(\phi_{ij}(\cdot))$ is learned and context-dependent (from $q$)
These distinctions highlight that while all three methods employ sampling for scalability, they differ fundamentally in whether the sampling distribution is learned, what structural elements are sampled, and where adaptivity occurs in the system architecture.
Appendix I Limitations
While BayesG demonstrates strong empirical performance and scalability on traffic signal control benchmarks, several limitations remain:
  • Fixed physical topology. Our approach assumes a predefined, static environment graph that restricts the set of possible neighbor interactions. While this models many real-world systems such as traffic networks, it may not generalize to domains with dynamic or learned topology [43, 44, 45].
  • Local observability. Each agent infers its latent communication graph based solely on local observations within its ego-graph. This limits the model’s ability to reason over long-range dependencies that may require more global context [46, 47].
  • Hyperparameter sensitivity. The performance of the learned mask depends on the Gumbel-softmax temperature and sparsity prior. Improper tuning may lead to either under- or over-pruning of edges, potentially degrading coordination [48, 49].
  • Computational cost. Compared to non-communicative methods, BayesG introduces additional overhead due to variational sampling and masked GNN computation. While still efficient in practice, training cost may increase with neighborhood size or message-passing depth.
We believe these limitations open avenues for future work on dynamic graph adaptation, broader task domains, and further efficiency improvements.